OpenStudy (anonymous):

two of the top grossing concert tours were by a jazz band and rock band. together the two tours visited 178 cities. the jazz band visited 92 cities more than the rock band how many cities did the jazz band visit? how many cities did the rock band visit?

OpenStudy (anonymous):

let J = jazz band let R = rock band R+J=178 J-R=92

OpenStudy (anonymous):

solve using substitution J = 92+R R+(92+R)=178 solve for J and R

OpenStudy (anonymous):

best best thank u veryy muchh